OpenStudy (anonymous):

Find the solution of this system of equations.. 3x = 9 - 6y 3x + 8y = 9 @babygirl11271963

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

this one youll want to use substitution for

OpenStudy (anonymous):

go to mathway it answers ur math questions

OpenStudy (anonymous):

you see 3x=9-6y. divide everything by 3 in that equation. you come to x=3-2y. no go to the second equation. SUBSTITUTE the first equation for x. 3(3-2y) +8y = 9. simplify it to 9-6y+8y=9. now put like terms together. 9+2y=9. subtract 9 from both sides. 2y=0 so y=0. now substitute 0 in for y in the first equation. 3x=9-6(0) so you have 3x=9. divide both sides by 3 and you get x=3. (3,0)

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

Thanks :D.

OpenStudy (anonymous):

welcome :)